OpenStudy (anonymous):

what does the term i^2 simplify?

OpenStudy (saifoo.khan):

-1

OpenStudy (anonymous):

i = sqrt -1 so square of this = -1

OpenStudy (anonymous):

i is the irrational \[\sqrt{-1}\] so i^2 is -1

OpenStudy (anonymous):

i^2 = -1 is a definition for i